# Enhancing Accuracy in Differentially Private Distributed Optimization Through Sensitivity Reduction

### v1

@@ Abstract
In this paper, we investigate the problem of differentially private distributed optimization. Recognizing that lower sensitivity leads to higher accuracy, we analyze the key factors influencing the sensitivity of differentially private distributed algorithms. Building on these insights, we propose a novel differentially private distributed algorithm that enhances optimization accuracy by reducing sensitivity. To ensure practical applicability, we derive a closed-form expression for the noise parameter as a function of the privacy budget. Furthermore, we rigorously prove that the proposed algorithm can achieve arbitrarily rigorous $\epsilon$-differential privacy, establish its convergence in the mean square sense, and provide an upper bound on its optimization accuracy. Finally, extensive comparisons with various privacy-preserving methods validate the effectiveness of our algorithm.
@@ 1 Introduction
In general, privacy-preserving distributed optimization algorithms can be categorized into three types: correlated randomness-based methods, encryption-based methods, and differential privacy-based methods. Correlated randomness-based methods achieve privacy protection by introducing temporally or spatially correlated randomness into optimization parameters [4, 5, 6, 7]. For example, [6] preserves the privacy of agents’ gradients by adding randomness to the stepsize and coupling weights during the initial iterations. By leveraging the robustness of dynamics, these methods can attain exact convergence. However, they necessitate an additional topological assumption to ensure privacy protection and cannot defend against external eavesdroppers. In contrast, encryption-based methods can effectively defend against external eavesdroppers by encrypting the transmitted information [8, 9, 10]. For example, in [10], agents use the Advanced Encryption Standard (AES) to encrypt transmitted information before sharing it with neighboring agents. However, encryption and decryption operations substantially increase the algorithm’s computational and communication overhead.
@@
Due to the wide applicability, simplicity, and strong theoretical privacy guarantees, Differential Privacy (DP) has emerged as the de facto standard for privacy protection in recent years [11]. Several studies have integrated DP into distributed optimization by injecting random noise into the information transmitted between agents [12, 13, 14, 15, 16]. For example, [12] proposes a differentially private algorithm based on distributed projected gradient descent, under the assumption that the gradients of all individual cost functions are bounded. However, since distributed gradient descent relies on inexact gradients, its convergence rate remains sublinear, even without the addition of DP noise. In contrast, the gradient tracking algorithm can achieve a linear convergence rate with a constant stepsize. To this end, [13] presents a differentially private algorithm based on gradient tracking that achieves linear convergence. However, this approach assumes that different objective functions in two adjacent problems share the same Lipschitz constant and convexity parameter. On the other hand, both [12] and [13] reveal the inherent trade-off between privacy and accuracy in differentially private algorithms. To overcome this challenge, [14] proposes a distributed algorithm based on gradient tracking that ensures both DP and optimality by incorporating decaying weakening factors. However, to achieve a finite privacy budget as the number of iterations approaches infinity, this method requires the assumption that different objective functions in adjacent problems have consistent gradients within the neighborhood of the optimal solution of one problem. In addition, in the presence of persistent DP noise, gradient tracking techniques suffer from noise accumulation, which negatively impacts the algorithm’s optimization accuracy. To mitigate this issue, [15] proposes a differentially private distributed algorithm based on robust gradient tracking. However, its applicability is restricted to undirected graphs due to the use of doubly stochastic weight matrices. To address this limitation, [16] employs a row and a column stochastic weight matrices to develop a robust gradient tracking-based differentially private algorithm that extends applicability to directed graphs. [15] and [16] also reveal the inherent trade-off between privacy and accuracy in differentially private algorithms. Furthermore, [16] relaxes the assumption that the gradients of all individual cost functions are bounded, instead assuming that the distance between the gradients of different objective functions in adjacent problems is bounded.
Despite significant advancements in differentially private distributed optimization, several critical challenges remain. Firstly, existing differentially private algorithms lack design principles to enhance their accuracy, as well as theoretical validation for such improvements. Secondly, there is no direct analytical relationship between the noise parameter and the privacy budget. As a result, it becomes impractical to determine the required noise level for a given privacy budget before the algorithm’s execution, thus limiting real-world deployment. Thirdly, existing convergence analysis techniques become inapplicable in the presence of noise injection and diminishing stepsize. To address the limitations, this paper proposes corresponding solutions, which, along with the contributions of this work, are summarized as follows:
@@
1) Enhancement of the accuracy of the differentially private distributed algorithm: We first reveal the relationship between the algorithm’s accuracy and sensitivity, demonstrating that lower sensitivity results in higher accuracy under the same privacy budget. This establishes a theoretical foundation for the design of the algorithm. Subsequently, we analyze the main factors affecting the sensitivity of the algorithm (Proposition 1). Building on this analysis, we propose an implicit gradient tracking-based differentially private distributed algorithm with lower sensitivity. Furthermore, we theoretically and empirically validate that the proposed algorithm achieves higher accuracy compared to existing methods [13, 15, 14, 16] under the same privacy budget.
2) Establishment of a closed-form expression for the noise parameter as a function of the privacy budget and reduction of the communication and computational burdens: For the proposed differentially private distributed algorithm, we derive a closed-form expression for the noise parameter as a function of the privacy budget. Based on this expression, it becomes feasible to directly compute the required noise parameter for any given privacy budget. Additionally, compared to [13, 15, 14, 16], the proposed algorithm reduces the number of communication variables by half, alleviating the communication burden. The number of injected noise is also halved, thereby reducing the computational burden of the algorithm.
@@
3) Theoretical analysis of $\epsilon$-DP and convergence: Under a weak assumption on the gradients of different objective functions in adjacent problems, we prove that the proposed algorithm with appropriate parameters can achieve rigorous $\epsilon$-DP (Theorem 4.5). Additionally, under the assumption that the objective function is Lipschitz smooth and strongly convex, we establish the convergence of the proposed algorithm with diminishing stepsize and noise parameter, as well as an upper bound for its optimization accuracy (Theorem 5.17). The theoretical results reveal the trade-off between the algorithm’s privacy and accuracy. Based on these results, we further provide two tuning strategies for the algorithm’s parameters.
@@
The remainder of this paper is organized as follows. Section 2 presents the problem formulation and provides relevant definitions of DP. Section 3 proposes a differentially private distributed algorithm with lower sensitivity. Sections 4 and 5 analyze the differential privacy guarantees and convergence of the proposed algorithm, respectively. Section 6 validates the effectiveness of the proposed algorithm through numerical simulations. Finally, Section 7 concludes the paper.
Notation: For any vector $\bm{v}\in\mathbb{R}^{p}$ and $1\leq a\leq\infty$, $\|\bm{v}\|_{a}$ represents the standard $L^{a}$-norm for $\bm{v}$. Without a subscript, $\|\bm{v}\|$ denotes $L^{2}$-norm of $\bm{v}$. For any matrix $M\in\mathbb{R}^{n\times p}$, $\|M\|_{a}$ represents the standard induced $L^{a}$-norm for $M$ such that $\|M\|_{a}\triangleq\sup_{\|x\|_{a}\not=0}\frac{\|Mx\|_{a}}{\|x\|_{a}}$. Without a subscript, $\|M\|$ denotes the Frobenius norm of $M$. $\bm{0},\bm{1}\in\mathbb{R}^{n}$ represent the vectors of all zeros and ones, respectively. $I\in\mathbb{R}^{n\times n}$ denotes the identity matrix. For a square matrix $M\in\mathbb{R}^{n\times n}$, $\rho(M)$ represents its spectral radius. Given a sequence of matrices $\{M_{k}\}$ with $M_{k}\in\mathbb{R}^{n\times n}$ and for $K_{1},K_{2}\geq 1$,
$\prod_{k=K_{1}}^{K_{2}}M_{k}=\left\{\begin{aligned} &M_{K_{2}}\cdots M_{K_{1}},&\text{if}\ K_{2}>K_{1},\\ &M_{K_{1}},&\text{if}\ K_{2}=K_{1},\\ &I,&\text{if}\ K_{2}<K_{1}.\end{aligned}\right.$
For the random variables $X$ and $Y$, $\mathbb{P}[X]$ represents the probability of $X$, $\mathbb{E}[X]$ denotes the expectation of $X$, and $\mathbb{E}[X|Y]$ is the conditional expectation of $X$ given $Y$. Given a constant $c>0$, $Lap(c)$ represents the Laplace distribution with the probability density function $p_{c}(x)\triangleq\frac{1}{2c}e^{-\frac{|x|}{c}}$. This distribution has a mean of zero and a variance of $2c^{2}$. For any $x,y\in\mathbb{R}$, it can be shown that $\frac{p_{c}(x)}{p_{c}(y)}\leq e^{\frac{|y-x|}{c}}$. If $x\sim Lap(c)$, $\mathbb{E}[x]=0$, $\mathbb{E}[|x|]=c$, and $\mathbb{E}[x^{2}]=2c^{2}$. For any matrix $M\in\mathbb{R}^{n\times p}$, $M\sim Lap(c)$ implies that each component $M_{ij}\sim Lap(c)$.
@@ 2 Problem Formulation
where $\bm{x}\in\mathbb{R}^{p}$ is the global decision variable and $f:\mathbb{R}^{p}\to\mathbb{R}$ is the average of all individual cost functions. Let $\bm{x}_{i}$ be a local copy of the decision variable for agent $i$. Denote the gradient of $f_{i}$ with respect to $\bm{x}_{i}$ as $\nabla f_{i}(\bm{x}_{i})$. For convenience, we further define $\mathbf{x}\triangleq\left[\bm{x}_{1},\cdots,\bm{x}_{n}\right]^{T}\in\mathbb{R}^{n\times p}$ and $\nabla F(\mathbf{x})\triangleq\left[\nabla f_{1}(\bm{x}_{1}),\cdots,\nabla f_{n}(\bm{x}_{n})\right]^{T}\in\mathbb{R}^{n\times p}$.
@@
Next, we introduce some notations. We describe the distributed optimization problem $\mathcal{P}$ in (1) using the tuple $(\mathcal{X},\mathcal{F},f,\mathcal{G}_{W})$, where $\mathcal{X}=\mathbb{R}^{p}$ is the domain of optimization. $\mathcal{F}\subseteq\{\mathcal{X}\to\mathbb{R}\}$ is a set of real-valued cost functions. $f(\bm{x})\triangleq\frac{1}{n}\sum_{i=1}^{n}f_{i}(\bm{x})$ with $f_{i}\in\mathcal{F}$ for each $i\in\mathcal{N}$. $\mathcal{G}_{W}$ is a communication graph induced by the weight matrix $W$. We represent a distributed optimization algorithm for solving the problem $\mathcal{P}$ by a mapping $Alg\left(\mathcal{P},\mathbf{x}(0)\right):\{\mathcal{P},\mathbf{x}(0)\}\to\mathcal{Z}$, where $\mathbf{x}(0)\in\mathcal{X}^{n}$ is the initial state of all agents. $\mathcal{Z}=\{\mathbf{z}(0),\mathbf{z}(1),\cdots\}\in\mathbb{O}$ is the sequence of shared information of all agents, also known as the observation sequence, and $\mathbb{O}$ denotes the set of all possible observation sequences. Furthermore, under a distributed optimization algorithm $Alg$, for $\mathcal{P}$, $\mathbf{x}(0)$, and $\mathcal{Z}$, we denote the algorithm’s internal states as $Sta(\mathcal{P},\mathbf{x}(0),\mathcal{Z})$.
@@
(Adjacency) Two distributed optimization problems $\mathcal{P}$ and $\mathcal{P}^{\prime}$ are adjacent, if the following hold:
@@
Regarding the condition on different cost functions in adjacent problems, [12] and [15] require that the gradients of all cost functions be bounded, i.e., for any $\bm{x}\in\mathcal{X}$ and $f_{i}\in\mathcal{F}$, $\|\nabla f_{i}(\bm{x})\|\leq\delta_{1}$ with $\delta_{1}>0$. In contrast, our requirement in Definition 1 is more relaxed, allowing for a broader class of cost functions. For example, the functions $f_{i_{0}}(\bm{x})=\frac{1}{2}\bm{x}^{T}\bm{x}$ and $f_{i_{0}}^{\prime}(\bm{x})=\frac{1}{2}\bm{x}^{T}\bm{x}+\frac{\delta}{p}\bm{x}$ satisfy our requirement, but their gradients are unbounded on $\mathbb{R}^{p}$. Furthermore, [13] assumes that $f_{i_{0}}$ and $f_{i_{0}}^{\prime}$ share the same Lipschitz and convexity constants. [14] assumes that $f_{i_{0}}$ and $f_{i_{0}}^{\prime}$ have identical gradients in the neighborhood of the solution of $\mathcal{P}$. These assumptions impose significantly stricter constraints compared to ours.

### v2

@@ Abstract
In this paper, we investigate the problem of differentially private distributed optimization. Recognizing that lower sensitivity leads to higher accuracy, we analyze the key factors influencing the sensitivity of differentially private distributed algorithms. Building on these insights, we propose a novel differentially private distributed algorithm for undirected graphs that enhances optimization accuracy by reducing sensitivity. To ensure practical applicability, we derive an explicit closed-form expression for the noise parameter as a function of the privacy budget. Moreover, we rigorously prove that the proposed algorithm can achieve arbitrarily rigorous $\epsilon$-differential privacy, establish its convergence in the mean square sense, and provide an upper bound on its optimization accuracy. Finally, extensive comparisons with various privacy-preserving methods validate the effectiveness of our algorithm.
@@ 1 Introduction
In general, privacy-preserving distributed optimization algorithms can be categorized into three types: correlated randomness-based methods, encryption-based methods, and differential privacy-based methods.
Correlated randomness-based methods achieve privacy protection by introducing temporally or spatially correlated randomness into the optimization process. [4] and [5] protect agents’ private information by injecting structured random noise adhering to a zero-sum structure into the interactive messages between agents. [6] further strengthens privacy by simultaneously introducing functional and state perturbations that also satisfy a zero-sum constraint. [7] safeguards agents’ privacy by injecting random noise into the dual variables directed towards a non-convergent subspace. [8] protects the privacy of agents’ gradients by adding randomness to the stepsizes and coupling weights during the initial iterations. [9] and [10], on the other hand, utilize state decomposition techniques to split the local state into public and private substates, transmitting only the substate unrelated to private information during communication, thereby achieving privacy preservation. Leveraging the robustness of the system dynamics, these methods can attain exact convergence. However, they require an additional topological assumption to ensure privacy protection and cannot defend against external eavesdroppers.
In contrast, encryption-based methods can effectively defend against external eavesdroppers by encrypting the transmitted information. [11] and [12] propose two privacy-preserving distributed optimization algorithms by integrating the Alternating Direction Method of Multipliers (ADMM) and Distributed Gradient Descent (DGD) methods with Homomorphic Encryption (HE). Considering the significant computational and communication overhead associated with HE, several efficiency optimization techniques can be adopted to mitigate these burdens. [13] combines quantization techniques with HE, significantly reducing the size of data to be encrypted and transmitted by converting high-precision gradients into low-bit integers. [14] integrates packing (batching) techniques with HE, reducing communication frequency by aggregating multiple data elements into a single plaintext or ciphertext before encryption and transmission. Moreover, replacing HE with lightweight cryptographic primitives, such as the Advanced Encryption Standard (AES) in [15] and functional encryption in [16], can reduce the overall computational and communication burden. Despite these advances, compared to other types of privacy-preserving methods, the encryption and decryption operations in encryption-based approaches still introduce additional computational or communication overhead.
@@
Due to the wide applicability, simplicity, and strong theoretical privacy guarantees, Differential Privacy (DP) has emerged as the de facto standard for privacy protection in recent years [17]. Several studies have integrated DP into distributed optimization by injecting random noise into the information transmitted between agents. [18] proposes a differentially private algorithm based on DGD, under the assumption that the gradients of all individual cost functions are bounded. However, since DGD relies on inexact gradients, its convergence rate remains sublinear, even without the addition of DP noise. In contrast, the Gradient Tracking (GT) algorithms can achieve linear convergence with a constant stepsize [19]. To this end, [20] presents a differentially private GT algorithm that achieves linear convergence over undirected graphs. This result is later extended to time-varying directed graphs by [21]. Both works, however, rely on the assumption that different objective functions in two adjacent problems share the same Lipschitz constant and convexity parameter. On the other hand, [18], [20], and [21] all reveal the inherent trade-off between privacy and accuracy in differentially private algorithms.
To optimize the privacy-accuracy trade-off in differentially private algorithms, various techniques have been proposed. [22] presents a GT-based distributed algorithm that ensures both DP and optimality by incorporating decaying weakening factors into weight matrices. However, to achieve rigorous DP, this method requires the assumption that adjacent objective functions have consistent gradients within the neighborhood of the optimal solution. In addition, in the presence of persistent DP noise, GT techniques suffer from noise accumulation, which negatively impacts the algorithm’s accuracy. To mitigate this issue, [23] proposes a differentially private distributed algorithm based on Robust Gradient Tracking (RGT) for undirected graphs. [24] later extends this approach to directed graphs. [25] proposes a distributed algorithm that integrates RGT with decaying weakening factors, ensuring almost sure convergence in the presence of information-sharing noise. As an enhancement, [26] further incorporates a variance reduction technique into this framework, which reduces the estimation error variance and relaxes the assumptions on network connectivity. However, whether these two methods can provide rigorous DP guarantees remains unclear.
Apart from the aforementioned research focused on suppressing or reducing the impact of noise, another line of work aims to optimize the noise itself, with the core idea being to dynamically adjust the scale and distribution of noise based on the characteristics of the model or data. The adaptive noise mechanism is a typical representative of this direction. For instance, in the federated learning setting, [27] proposes a method combining adaptive gradient clipping and adaptive noise scaling, which better aligns the noise with the training phase and client heterogeneity. Furthermore, [28] introduces a component-wise adaptive noise mechanism that takes into account the heterogeneity in parameter magnitudes. By estimating the sensitivity of each parameter component and adding Gaussian noise scaled accordingly, this approach prevents excessive distortion to parameters with smaller magnitudes. On the other hand, from the perspective of system and communication architecture, introducing the shuffle model to achieve “privacy amplification” has emerged as another important approach to enhance privacy. This model incorporates an anonymization step: users first perturb their data locally, after which a trusted shuffler randomly permutes all user messages before sending them to the analyzer. This process severs the association between outputs and their senders, thereby providing stronger central DP guarantees for aggregated results compared to pure local perturbation [29, 30, 31]. However, how to apply adaptive noise mechanisms and privacy amplification strategies to general differentially private decentralized optimization remains an open problem requiring in-depth research.
Despite significant advancements in differentially private distributed optimization, several critical challenges remain. Firstly, existing differentially private algorithms lack clear design principles to systematically improve accuracy, as well as theoretical frameworks to validate such improvements. As a result, efforts to enhance accuracy often rely on heuristic tuning rather than guided optimization, and it remains challenging to rigorously discern whether a modification truly improves performance or merely shifts the trade-off elsewhere. Secondly, while relationships between privacy and noise have been studied, there remains a lack of an explicit, closed-form expression that can directly translate a desired privacy budget into actionable noise parameters prior to execution. This makes it difficult to configure algorithms predictably and limits their practical deployment. Thirdly, the use of diminishing stepsizes introduces a fundamental analytical hurdle: the resulting state transition matrix of the algorithm no longer has a spectral radius strictly less than 1. This, combined with persistent noise injection, invalidates standard convergence analysis techniques, making it difficult to establish rigorous performance guarantees. To address these challenges, this paper proposes corresponding solutions, which, along with the contributions of this work, are summarized as follows:
@@
1) A Systematic, Sensitivity-Reduced Design for Accuracy Enhancement. To overcome the lack of design principles, we first reveal that algorithm sensitivity is the key determinant of accuracy under a fixed privacy budget. This novel perspective enables the systematic derivation of design principles for sensitivity reduction (Proposition 1). Implementing these principles, we propose an implicit gradient tracking-based differentially private distributed algorithm for undirected graphs that inherently achieves lower sensitivity. Both theoretical analysis and experiments confirm that the proposed algorithm achieves higher accuracy under the same privacy budget compared to existing methods in [20, 21, 23, 22, 24].
2) An Explicit, Direct Noise Calibration Rule and Halved Overhead. To address the impracticality of implicit noise-privacy relationships, we derive an explicit, directly usable closed-form expression for the noise parameter as a function of the privacy budget. This expression serves as a direct design tool, enabling the exact computation of noise parameters from any given privacy budget before execution. Additionally, the proposed algorithm halves both the number of communication variables and noise injections compared to existing methods in [20, 21, 23, 22, 24], significantly reducing communication and computational overhead.
@@
3) Theoretical Analysis of $\epsilon$-DP and Convergence. We prove that the proposed algorithm with appropriate parameters can achieve rigorous $\epsilon$-DP under a weak gradient dissimilarity condition (Theorem 4.5). Furthermore, to address the analytical difficulty arising from diminishing stepsizes and persistent noise, specifically, the resulting non-contractive dynamics, we develop an analysis framework to establish mean-square convergence and derive an upper bound on the optimization accuracy (Theorem 5.18). This bound explicitly quantifies the privacy-accuracy trade-off. Based on the theoretical results, we provide two practical parameter tuning strategies.
@@
The remainder of this paper is organized as follows. Section 2 presents the problem formulation and provides relevant definitions of DP. Section 3 proposes a differentially private distributed algorithm with lower sensitivity. Sections 4 and 5 analyze the DP guarantees and convergence of the proposed algorithm, respectively. Section 6 validates the effectiveness of the proposed algorithm through numerical simulations. Finally, Section 7 concludes the paper.
Notation: Table 1 lists the key notations used in this paper; others are defined in their respective sections.
Table 1: Summary of Key Notations
Symbol Meaning
A. Key Problem and Algorithmic Parameters
$n$ Number of agents
$p$ Dimension of the decision variable
$_{i}, ff$ Local and global cost functions
$\bm{x}_{i}$, $\mathbf{x}$ Local variable of agent $i$ and stack matrix of all $\bm{x}_{i}$
$\bm{x}^{*}$ Optimal solution to the global problem
$L$, $\mu$ Lipschitz smoothness and strong convexity constants
$\bm{y}_{i}$ Auxiliary variable for agent $i$ (not communicated)
$\bm{z}_{i}$ Noisy state of agent $i$ (shared with neighbors)
$\bm{\xi}_{i}$ Laplace noise added by agent $i$
$\alpha_{k}$ Stepsize at iteration $k$
$\beta$ Gain parameter for the auxiliary variable
$\nu_{k}$ Parameter of the Laplace noise at iteration $k$
$\epsilon$ Privacy budget
$\delta$ Bound on gradient differences for adjacent problems
$\Delta(k)$ Sensitivity of the algorithm at iteration $k$
B. General Mathematical Symbols
$\|\bm{v}\|_{a}$ Standard $L^{a}$-norm of vector $\bm{v}$
$\|\bm{v}\|$ $L^{2}$-norm of vector $\bm{v}$
$\|M\|_{a}$ Standard induced $L^{a}$-norm of matrix $M$
$\|M\|$ Frobenius norm of matrix $M$
$\mathbf{0}$, $\mathbf{1}$ All-zeros and all-ones vectors
II Identity matrix
$\rho(M)$ Spectral radius of square matrix $M$
$\mathbb{P}[\cdot]$, $\mathbb{E}[\cdot]$ Probability and expectation operators
$Lap(c)$ Laplace distribution with scale parameter $c$
@@ 2 Problem Formulation
where $\bm{x}\in\mathbb{R}^{p}$ is the global decision variable and $f:\mathbb{R}^{p}\to\mathbb{R}$ is the average of all cost functions. This paper aims to protect the privacy of each agent’s local cost function $f_{i}$, which constitutes sensitive information. We consider an adversary with arbitrary computational power that can observe all shared information among all agents. This adversary may be an external eavesdropper or any coalition of internal honest-but-curious agents, with knowledge of the network topology, the algorithm’s parameters and rules, and the cost functions of all non-target agents. The privacy objective is to prevent this adversary from inferring the private cost function $f_{i}$ of any individual agent. Through DP, we ensure the adversary cannot reliably determine the identity or exact form of any agent’s cost function from the algorithm’s outputs. Let $\bm{x}_{i}$ be a local copy of the decision variable for agent $i$. Denote the gradient of $f_{i}$ with respect to $\bm{x}_{i}$ as $\nabla f_{i}(\bm{x}_{i})$. For convenience, we further define $\mathbf{x}\triangleq\left[\bm{x}_{1},\cdots,\bm{x}_{n}\right]^{T}\in\mathbb{R}^{n\times p}$ and $\nabla F(\mathbf{x})\triangleq\left[\nabla f_{1}(\bm{x}_{1}),\cdots,\nabla f_{n}(\bm{x}_{n})\right]^{T}\in\mathbb{R}^{n\times p}$.
@@
Remark 1
(On the network topology) Assumption 2 requires the communication graph to be undirected and connected, and the weight matrix to be doubly-stochastic. While this is more restrictive than the directed-graph settings considered in some existing privacy-preserving algorithms (e.g., [22, 24, 21]), it is adopted here to enable a clear investigation of the fundamental challenges in differentially private distributed optimization. Extending the design methodology and theoretical results of this work to directed graphs, which typically rely only on row- or column-stochastic weights, is an important and nontrivial direction for future research.
Next, we introduce some notations. We describe the distributed optimization problem $\mathcal{P}$ in (1) using the tuple $(\mathcal{X},\mathcal{F},f,\mathcal{G}_{W})$, where $\mathcal{X}=\mathbb{R}^{p}$ is the domain of optimization. $\mathcal{F}\subseteq\{\mathcal{X}\to\mathbb{R}\}$ is a set of real-valued cost functions. $f(\bm{x})\triangleq\frac{1}{n}\sum_{i=1}^{n}f_{i}(\bm{x})$ with $f_{i}\in\mathcal{F}$ for each $i\in\mathcal{N}$. $\mathcal{G}_{W}$ is a communication graph induced by the weight matrix $W$. We represent a distributed optimization algorithm for solving the problem $\mathcal{P}$ by a mapping $Alg\left(\mathcal{P},\mathbf{x}(0)\right):\{\mathcal{P},\mathbf{x}(0)\}\to\mathcal{Z}$, where $\mathbf{x}(0)\in\mathcal{X}^{n}$ is the initial state of all agents. $\mathcal{Z}=\{\mathbf{z}(0),\mathbf{z}(1),\cdots\}\in\mathbb{O}$ is the sequence of shared information of all agents, also known as the observation sequence, and $\mathbb{O}$ denotes the set of all possible observation sequences. Note that since each $\mathbf{z}(k)$ takes values in the continuous space $\mathcal{X}^{n}$, the observation set $\mathbb{O}$ is continuous. Furthermore, under a distributed optimization algorithm $Alg$, for $\mathcal{P}$, $\mathbf{x}(0)$, and $\mathcal{Z}$, we denote the algorithm’s internal states as $Sta(\mathcal{P},\mathbf{x}(0),\mathcal{Z})$, which encompass all data and variables that are stored locally by each agent and not directly shared with other agents.
@@
(Adjacency [24]) Two distributed optimization problems $\mathcal{P}$ and $\mathcal{P}^{\prime}$ are adjacent, if the following hold:
@@
Regarding the condition on different cost functions in adjacent problems, [18] and [23] require that the gradients of all cost functions be bounded, i.e., for any $\bm{x}\in\mathcal{X}$ and $f_{i}\in\mathcal{F}$, $\|\nabla f_{i}(\bm{x})\|\leq\delta_{1}$ with $\delta_{1}>0$. In contrast, our requirement in Definition 1 is more relaxed, allowing for a broader class of cost functions. For example, the functions $f_{i_{0}}(\bm{x})=\frac{1}{2}\bm{x}^{T}\bm{x}$ and $f_{i_{0}}^{\prime}(\bm{x})=\frac{1}{2}\bm{x}^{T}\bm{x}+\frac{\delta}{p}\bm{x}$ satisfy our requirement, but their gradients are unbounded on $\mathbb{R}^{p}$. Furthermore, [20] and [21] assume that $f_{i_{0}}$ and $f_{i_{0}}^{\prime}$ share the same Lipschitz and convexity constants. [22] assumes that $f_{i_{0}}$ and $f_{i_{0}}^{\prime}$ have identical gradients in the neighborhood of the solution of $\mathcal{P}$. These assumptions impose significantly stricter constraints compared to ours.